### v1

# It’s a (Blind) Match! Towards Vision-Language Correspondence without Parallel Data

Abstract
The platonic representation hypothesis suggests that vision and language embeddings become more homogeneous as model and dataset sizes increase. In particular, pairwise distances within each modality become more similar. This suggests that as foundation models mature, it may become possible to match vision and language embeddings in a fully unsupervised fashion, i.e. without parallel data. We present the first feasibility study, and investigate conformity of existing vision and language foundation models in the context of unsupervised, or “blind”, matching. First, we formulate unsupervised matching as a quadratic assignment problem and introduce a novel heuristic that outperforms previous solvers. We also develop a technique to find optimal matching problems, for which a non-trivial match is very likely. Second, we conduct an extensive study deploying a range of vision and language models on four datasets. Our analysis reveals that for many problem instances, vision and language representations can be indeed matched without supervision. This finding opens up the exciting possibility of embedding semantic knowledge into other modalities virtually annotation-free. As a proof of concept, we showcase an unsupervised classifier, which achieves non-trivial classification accuracy without any image-text annotation.
1 Introduction
Vision and language foundation models are changing the research landscape in computer vision and natural language processing. On the one hand, computer vision has widely adopted deep image representations, such as DINO [3], across virtually all applications, including semantic understanding [15] and 3D reconstruction [50]. On the other hand, large language models (LLMs) [e.g., 10] are already excelling at imitating natural language, performing comparably or better than humans in language translation and basic reasoning [e.g., see 53, for a survey]. Yet, we gravely lack in understanding the conformity between these models. Previous work offers some analysis under the lens of a single modality or in the cross-modal scenario with focus on label efficiency [35, 34]. In particular, the latter body of work pursued dense alignment between vision and language embeddings from a few correspondences. Taking this scenario to the extreme, we study “blind” matching, illustrated in Fig. 1. Specifically, we aim to find vision-language correspondence without any parallel data.
Vision and language representations may seem incompatible at first. Indeed, there are significant differences in the training process, data distribution, and model architectures. However, the platonic representation hypothesis [17] posits that textual and visual representations of the same concepts converge to geometrically similar latent spaces. It asserts that models trained on large, diverse datasets encounter the same phenomena and, more importantly, the same relationships between the phenomena. As Fig. 1 conceptually illustrates, existing vision and language models already exhibit a high degree of vision-language compatibility. As Huh et al. [17] also show, the alignment grows with larger sizes of models and training corpora.
Capitalizing on the platonic representation hypothesis, we study vision-language alignment from a completely unsupervised perspective. Developing techniques towards such an analysis is valuable for two reasons. First, it will provide us with tools to study the conformity between vision and language models also on the abundance of unlabelled data. Second, establishing correspondence between semantic concepts (encapsulated by text) and images paves the way for purely unsupervised visual recognition. In fact, our study culminates in an intriguing proof-of-concept application, unsupervised classifiers, which perform (semantic) image recognition without any paired data.
Contributions. As our technical contribution to the study of blind vision-language alignment, we first formulate the task as a quadratic assignment problem (QAP). This formulation requires only pairwise distances within each modality as the input. We then discuss and analyze the metrics behind the pairwise relationships. Moving forward, we revisit and extend a solver for QAPs, exploiting a specialized heuristic towards memory-efficient and accurate vision-language matching. We thoroughly compare our solver extension with previous heuristics for QAPs and vision-language matching, establishing that our solver leads to better optima and provides tighter bounds. Lastly, we define optimal criteria for finding suitable subsets of matching problems as a p-dispersion-sum problem [24].
As our main empirical contribution, we conduct a large-scale study involving a total of 33 vision and 27 language models on four datasets. Our study reveals that despite the complexity of the underlying optimization problem, we can establish non-trivial solutions in some of the vision-language assignment problems across varying problem sizes. Intriguingly, this finding enables unsupervised classification: classifying images into semantic categories without any parallel supervision.
2 Related work
Vision and language foundation models. Unsupervised learning has enabled large-scale pre-training, both in the vision [5, 3, 16, 36] and language domains [10, 43, 7]. While semantics are inherent to language, the emergence of semantics in unsupervised vision models is yet not well-understood [3]. Unlike unimodal foundation models, joint vision-text encoders, such as CLIP [40] and ALBEF [26], as well as generative text-to-image models [41, 45], leverage a shared vision-language embedding space [42].
Zero-shot stitching. Zero-shot stitching recovers dense alignment between two embedding spaces, such as vision and text. Typically, the task assumes a sparse set of available ground-truth image-text pairs. Stitching two unimodal pre-trained vision and language encoders is possible using relative representations [35, 34], or with the help of a geometric distortion penalty [20]. Maiorca et al. [29] estimate an affine transform enabling vision-text alignment of pre-trained modality-specific encoders. Similar to our work, Maniparambil et al. [30] study the similarity of the embedding spaces in the vision and text-based encoders. However, their evaluation does not consider the fully unsupervised alignment, which is the main focus of our study.
Gromov-Wasserstein distance. One possibility to formulate the assignment problem between two sets is the Gromov-Wasserstein distance, which computes pairwise distances within each set. The exact computation of the Gromov-Wasserstein distance is NP-hard; thus, its practical application has remained limited. The few examples include shape matching [32], graph matching and partitioning [51, 52], and point-cloud matching [37]. Alvarez-Melis and Jaakkola [1] apply the Gromov-Wasserstein distance for cross-lingual alignment without parallel data. This approach realizes downstream tasks, such as word translation, in an unsupervised fashion. Chen et al. [4] leverage the Gromov-Wasserstein distance for cross-modal alignment in model training. In contrast to these works, we perform alignment between language and vision models without parallel data and any parameter training.
QAP solvers. The quadratic assignment problem (QAP) is generally NP-hard [46]. However, efficient approximations exist for specific problems [38, 18]. There are many heuristics for general QAPs, which can be divided into primal and dual methods [14]. The primal methods try to solve an approximation of the original problem, e.g. by relaxing the constraints or limiting the search space. These include the fast approximate QAP algorithm (FAQ) [49], the 2-opt algorithm [8], and optimal transport Gromov-Wasserstein methods [38]. Specific to our task, LocalCKA [30] uses the centered kernel alignment (CKA) metric and approximates the QAP with a linear assignment problem for finding vision-language correspondences. Dual methods maximize the dual of the QAP, e.g. by using block coordinate ascent [13] or subgradient methods [47]. For our heuristic, we use a dual ascent algorithm from operations research, which we refer to as Hahn-Grant solver [13] (and detail further in Sec. 4.1). Finally, many methods, including ours, combine primal and dual solvers. For example, Hutschenreiter et al. [18] uses fusion moves in combination with proposals from the dual formulation of the linearized QAP, thereby iteratively improving the objective. This has been shown as a useful heuristic for many computer vision problems [14]. Commercial solvers for QAPs exist, and we mainly compare with Gurobi [12], a highly optimized and general solver for mixed integer programming.
3 Preliminaries
Here, we establish the link between the vision and language embeddings by formulating an alignment problem via pairwise similarities within each modality.
Let us find a correspondence between vision and language representations of classes. Given a set of images and text descriptions , we average their embeddings from a pre-trained vision model and a language model for each class :
Moreover, we -normalize each output of the model and the resulting averaged embeddings. We can then evaluate their similarity in each space individually using similarity measures and . The pairwise similarity matrices (or kernels) are then defined as
We measure the distortion of the pairwise similarities in vision and language spaces with a distance function . As its two arguments, the function takes the pairwise, possibly permuted, distance matrices from vision and language. Summing over all distortion elements yields the distortion between the vision and language embeddings:
Subject to a specific instantiation of , this formulation is general and can accommodate many existing distance (and similarity) measures. One notable example is the Gromov-Wasserstein distance [33]. Other examples include the mutual k-nearest neighbors (Mutual k-NN) [17] and the centered kernel alignment (CKA) [22]. Note that these are similarity, not distance measures. Without loss of generality, we will use the negative inner product with these measures and build our theory around minimizing a distance measure (see Appx. B).
We now demonstrate that the pairwise distortion is indicative of vision-language alignment. Following Maniparambil et al. [30], we begin with aligned matrices and , and gradually permute the language similarity matrix . For each permutation level, we randomly shuffle an increasing fraction of the language embeddings. We measure the alignment in terms of Mutual k-NN and define
where is the random permutation. Mutual k-NN measures the average overlap between the nearest neighbors in both modalities. In Fig. 2, we report the results for the CocoCaptions dataset [6], where we sample 100 random permutations for each permutation level. We observe that decreases strictly monotonically as the amount of shuffling increases. This observation implies that, on average, pairwise distances between two semantically related vision and language embeddings are more similar than between two unrelated embeddings. We confirmed this behavior for all considered datasets (e.g., ImageNet-100) and kernels (e.g., Gromov-Wasserstein distance).111The results also hold for randomly initialized vision and language networks. See Appx. C for discussion and further results.
4 Blind matching of vision and language
We now show that the above formulation of the vision-language alignment in terms of minimizing leads to a quadratic assignment problem (QAP). Then, we introduce the factorized Hahn-Grant solver, which efficiently finds an approximate solution to the QAP.
[…]
Here, is the group of permutation matrices, i.e. binary matrices in which the columns and the rows sum up to one:
4.1 The factorized Hahn-Grant solver
Solving the QAP in Eq. 6 by enumeration is not scalable, because there are different permutations. Furthermore, finding the global optimum for a general QAP is NP-hard [46]. In practice, even proprietary solvers like Gurobi [12] already fail for problem instances of size (see Sec. 5.3). Nevertheless, as we discussed in Sec. 2, one can recover the global solution or approximate it to a sufficient degree for some specific problem domains [49, 38]. Our main technical contribution is an extension of the Hahn-Grant solver that allows efficient matching of vision and language without any paired data.
[…]
The solver transfers the cost from the quadratic term to a linear and then a constant term by solving linear assignment problems (LAPs). The cost of the LAPs from the quadratic term can then be moved to leader (Lines 13-14), and the cost from the leader LAP can be moved to the constant term (Lines 6-7). On the other hand, the dual solutions of the LAPs are subtracted from the cost tensor and leader (Line 8 and Line 15). Moreover, in each step, both and are used to improve the dual (Lines 11-12). At the end of each step, the non-zero values of leader are evenly distributed back to the corresponding submatrices in (Line 9). Because of the redistribution, the next iteration can use these values to make further progress. The algorithm terminates once no improvement can be made. This means that the leader LAP can be solved with cost zero. Therefore, each of the LAPs corresponding to non-zero indices of the solution also has a zero-cost solution. If there is a permutation matrix that solves all of these LAPs, one has found the global optimum. In the algorithm, the constant maintains a lower bound of the original QAP. Since we solve LAPs in each iteration, the algorithm has a total runtime of and a memory complexity of .
The factorized Hahn-Grant solver. We adapt the Hahn-Grant solver for vision-language matching with three key extensions: (1) We also search for primal solutions, (2) reduce the memory complexity from to , and (3) use a faster solver for the LAPs.
We search for primal solutions in two stages. First, we initialize the problem with the FAQ [49] and 2opt [8] heuristics, using the best result from 100 random seeds (Line 4). Second, we evaluate the solutions of each LAP on the QAP (Line 7 and Line 14). This does not add considerable overhead, because the assignment was already calculated as a by-product. Therefore, we only need to evaluate the QAP objective functions once more. If the algorithm converges to a solution, that solution also needs to solve at least one of the LAPs [13], and our heuristic will find the global optimum. If the algorithm does not converge, we output the lowest-cost solution found during the run.
To reduce the memory requirements, we first observe that the commonly used distortion measures are decomposable as
[…]
This is an instance of a Koopmans-Beckmann QAP [21] with cost matrices and . Although these are only two matrices, applying the Hahn-Grant solver would require us to compute and store the full tensor . Instead of updating this tensor in-place, we store them separately in and . Therefore, the memory requirement is only , instead of .
Finally, we also experimented with the auction and the Jonker-Volgenant algorithm, instead of the Hungarian algorithm [25]. Despite the expected runtime of for the auction algorithm [2], we observed a faster convergence in all of our experiments using a specialized C++ implementation [31] of the Jonker-Volgenant algorithm [19] with a total asymptotic runtime of . Appx. D provides more details and a thorough comparison between our solver and the original Hahn-Grant solver.
4.2 Finding optimal matching problems
Despite the strong correlation between the distortion metric and the matching accuracy (cf. Fig. 2), the accuracy varies significantly depending on the choice of categories. This implies that a non-trivial matching for an arbitrary choice of semantic categories is not guaranteed. Nevertheless, we can find a subset of matching problems where a non-trivial match is very likely, even for large problem sizes ().
[…]
Therefore, we can find the subset that maximizes the alignment, i.e.
[…]
The non-zero entries of correspond to the classes in the subset . This problem is an instance of the p-dispersion-sum problem [24]. Although the problem is NP-hard in general [39], we can find reasonable approximations using Gurobi, as we demonstrate in Sec. 5.2.
5 Experiments
Our experiments encompass three studies and an example application. Our first study in Sec. 5.1 focuses on a small problem size (), where we can analyze the properties of the global solution in a tractable manner. We find that for an overwhelming number of pre-trained vision and language models, the optimal solutions w.r.t. the Gromov-Wasserstein distance correspond to non-trivial matching between the two modalities, well above the random chance. Sec. 5.2 presents the second study, where we employ our factorized Hahn-Grant solver to find solutions of larger-scale problems. This study reveals that there is a subset of larger matching problems where our solver finds a non-trivial match without any paired data. The third study in Sec. 5.3 compares our factorized Hahn-Grant solver to alternative optimization algorithms in terms of solution optimality. In particular, we observe that only the global optimum corresponds to non-trivial results. Finally, Sec. 5.4 extends these studies with an exciting application, introducing unsupervised classifiers.
Pre-trained models. In total, we consider vision models and language models using a variety of pre-trainings, architectures and sizes. For vision, we use self-supervised (DINO [3], DINOv2 [36]), fully-supervised (DeiT [48], ConvNeXt [27]), and vision-language supervised (CLIP [42]) models. For language, we use different models from the SentenceTransformers library [43] and CLIP [42]. Appx. E elaborates further.
Setup. Given an image classification dataset, we extract the vision embeddings for each image and compute the average vision embedding for each class. Similarly, we average the language embedding from multiple prompts to get a language representation. To observe the effect of small perturbations in the pairwise distances on the matching, we compute the vision representation using random subsets of the data. In our preliminary analysis, we found the Gromov-Wasserstein distance to be superior to CKA for matching vision and language.222See Appx. B for the comparison. Therefore, we use the Gromov-Wasserstein distance in all upcoming experiments.
5.1 Small-scale matching
Setup. For small matching problems (e.g. ), we can exactly enumerate the solution space, which offers an ideal testbed for our Gromov-Wasserstein measure. We perform matching on CIFAR-10 [23] and CINIC-10333CINIC-10 is a subset of ImageNet [44] with 10 classes from CIFAR-10. Our only adaption is that we keep the image resolution from ImageNet. [9].
Results. We show the matching accuracy for different vision models in Fig. 4. On the left, we use all-mpnet-base-v2 evaluated on CIFAR-10 and, on the right, All-Roberta-large-v1 on CINIC-10. More combinations are given in Sec. F.1. We observe that most models perform better than 10% accuracy, which is the baseline for random permutations. Moreover, the DINOv2 model can be matched with the highest accuracy of 72% on CIFAR-10 and 100% on CINIC-10. Curiously, we observe that it is the pre-training strategy that seems to have a larger impact than the model size. On average for all language models, the DINOv2 models are 5.2% better on CIFAR-10 and 7.7% better on CINIC-10 than the second-best pre-training strategy.
5.2 Larger-scale matching
We now investigate matching of larger problems (). For vision-language matching, a larger problem implies a finer granularity of the classes (apart from the computational considerations). This means that the pairwise distances must consistently encode more nuanced similarities.
Setup. We evaluate the matching accuracy on subsets of different sizes for ImageNet-100 [44] and CIFAR-100 [23]. Following Sec. 4.2, we hypothesize that not all classes are represented similarly. Therefore, we use our optimization problem from Sec. 4.2 to find subsets of classes that are well-aligned. We evaluate the alignment of the largest DINOv2, CLIP, and DeiT models with all-mpnet-base-v2.
Results Fig. 5 summarizes the results for this experiment for ImageNet-100 (Fig. 5, left) and for CIFAR-100 (Fig. 5, right). We observe that all models have high accuracy for small problem sizes, and that performance drops for larger problem sizes. Thus, for each model, there is a set of classes that is represented similarly. However, some categories adversely affect the matching, suggesting their different representation in vision and language models. As somewhat expected, CLIP outperforms the other two models in most cases, indicating that more classes share similar representation with the language model for CLIP than for DeiT and DINOv2. On CIFAR-100, the performance of DINOv2 declines more sharply than for the other models. Thus, some classes degrade the matching, if added. These results suggest that language supervision during training can align pairwise similarities, especially for fine-grained classes. Nevertheless, the self-supervised DINOv2 also learns aligned pairwise distances for some of the classes.
5.3 Solver comparison
We now show that previous solvers are often unable to obtain meaningful assignments for our matching problems.
Setup. We compare the solvers both in the small-scale setting from Sec. 5.1 and in the larger-scale setting using the optimal matching problems. We compare to the computer vision solver MPOpt [18], the commercial solver Gurobi [12], the fast approximate QAP algorithm (FAQ) [49], entropic Gromov-Wasserstein optimal transport (OT)444See Appx. G for how to use OT for QAPs. [38], and LocalCKA [30]. For both experiments, we use CLIP ViT-L/14@336px and all-mpnet-base-v2. More details and other settings are available in Sec. F.3.
Solver | Accuracy (%) | Cost | Global? (%) |
Random | ±10 | ±0.30 | ±0 |
LocalCKA [30] | ±9 | ±0.27 | ±0 |
OT [38] | ±4 | ±0.11 | ±0 |
FAQ [49] | ±3 | ±0.08 | ±0 |
MPOpt [18] | ±12 | ±0.02 | ±51 |
Gurobi [12] | ±0 | ±0.02 | ±0 |
Ours | ±0 | ±0.02 | ±0 |
Results. In the small-scale setting, we can compute the global optimum with enumeration. In Tab. 1, we compare the matching accuracy, the cost after optimization, and the percentage that the global optimum is reached for every method. For LocalCKA, the CKA metric is minimized using a linear approximation. We observe that minimizing this approximation does not lead to better results than random permutations, both in terms of Gromov-Wasserstein loss and the matching accuracy. Both local solvers (OT, FAQ) are able to minimize the loss, but are stuck in a local minimum, thus their accuracy is not better than the random baseline. This shows that local optima are not sufficient to find a useful match. MPOpt manages to approach the optimal objective, but only reaches it about half the time. In contrast, both our heuristic and Gurobi find the global optimum every time, resulting in accuracy of 60%.
Fig. 6 reports the results on the larger-scale problem. We show the cost and the lower bound for each problem size. For ease of comparison and for computational reasons, we select only one set of classes for each problem size. Consistent with the small-scale experiment, LocalCKA does not perform better than random and produces out-of-bound values. Interestingly, OT and FAQ are able to solve some instances with global optimality. We suspect that this is because the optimal problem may have a specialized structure. Therefore, the relaxations can lead to the global optimum for such problems. Already at size , all but our method fail to reach the global optimum. For problem sizes larger than 60, Gurobi does not produce a solution within the given time limit. Our method consistently provides the tightest bounds, provably leading to a global optimum up to size . Furthermore, it achieves the best primal solution for most sizes, even surpassing Gurobi on problems of size .
5.4 Application: Unsupervised classifiers
We introduce a proof-of-concept application exploiting our formulation of the unsupervised vision-language matching.
Setup. Assuming no annotation, we cluster image representations using K-Means [28]. We use the cluster centers for one-to-one matching with language embeddings using our QAP formulation, relying only on the pairwise distances. Given the solution, the elements of each cluster acquire the label of the corresponding language embedding.
Model | All-Roberta-large-v1 (%) | all-mpnet-base-v2 (%) |
CLIP | ±0.5 | ±0.4 |
DeiT | ±3.4 | ±7.5 |
DINOv2 | ±9.2 | ±0.4 |
Results. We evaluate our unsupervised classifier with several vision and language models in Tab. 2. The best accuracy is achieved by the DINOv2 model with 51.5% for All-Roberta-large-v1. However, all of the considered models achieve a higher accuracy than the random baseline of 10%. These results are clearly inferior to the supervised models. Nevertheless, to the best of our knowledge, this is the first instance of fully unsupervised classification. Furthermore, we show in Sec. F.2 that previous solvers are not able to find meaningful matches in this setting and analyze how clustering changes the matching performance.
6 Discussion
Can we match vision and language without paired data? Our small-scale experiments (cf. Sec. 5.1) indeed suggest that most vision and language models can be matched with high accuracy. This implies that on the tested datasets, the pairwise relations are similar. Furthermore, our larger-scale experiment (cf. Sec. 5.2) shows that it is also possible to match larger sets of fine-grained concepts. However, it also shows that some classes have different representation in vision and language, hence cannot be matched reliably.
Can we match arbitrary embeddings? Not with existing models yet. Our larger-scale experiments in Sec. 5.2 show that vision and language models may encode semantic concepts differently. This is not surprising, since some concepts are abstract and appear only in one of the modalities, e.g. “freedom of speech” [17]. For such concepts, conforming vision-language embedding spaces without explicit supervision are highly unlikely. Another issue is symmetry: vision-language matching may be ambiguous due to multiple (equivalent) local minima. Finally, the computational complexity scales in the order of . Therefore, our solver does not scale to problems of sizes larger than those studied in Sec. 5.2. Even for problems of size , the bound between the primal and the dual is significant. Thus, reaching the global optimum becomes infeasible.
Which vision model can be matched best? Our small-scale experiment in Sec. 5.1 indicates that DINOv2 models can be matched best. However, on CIFAR-100 in the large-scale experiment, we observe that CLIP and DeiT perform better, especially on larger problems. This suggests that DINOv2 consistently represents broader concepts, whereas CLIP and DeiT better encode fine-grained relationships.
7 Conclusion
In this work, we investigated the “blind” alignment of vision and language embedding spaces, i.e. without parallel data. Our contribution combines a technical and an empirical element. As a technical contribution, we showed that unsupervised alignment is an instance of a quadratic assignment problem, QAP. This problem formulation includes only the pairwise similarities in each embedding space. We proposed a new heuristic for solving the QAP, outperforming previous solvers on the alignment problem. On the empirical side, we extensively study the alignment of pre-trained vision and language models. Our analysis establishes feasibility of a non-trivial match between vision and language. Although solving general and large-scale matching problems remains an open challenge, the established feasibility creates exciting avenues for novel applications. The presented proof-of-concept unsupervised classifier is an intriguing first step in this direction.
Acknowledgments. We thank the anonymous reviewers for their valuable feedback. This work was supported by the Federal Ministry for the Environment, Nature Conservation, Nuclear Safety and Consumer Protection (BMUV) through the AuSeSol-AI project (grant 67KI21007A), and by the TUM Georg Nemetschek Institute Artificial Intelligence for the Built World (GNI) through the AICC project.
Supplementary Material
Appendix A Overview
Appx. B shows that our QAP formulation is general enough to accommodate the commonly used distance and similarity measures: Mutual k-NN, CKA and the Gromov-Wasserstein distance.
Appx. D compares the original Hahn-Grant solver to our factorized Hahn-Grant solver including a correctness proof, implementation details and an ablation showing the improvements in primal and dual estimates by our solver.
Appx. E gives more details on the used models, datasets and specific setups for the individual experiments.
Appx. F provides results for other model and dataset combinations. It shows that pre-training appears more important than the model size for the matching accuracy in small-scale alignment. Also, it shows that the results for solver comparison consistently hold for all considered model and dataset combinations including unsupervised classification.
Appx. G shows that the Gromov-Wasserstein optimal transport solvers can be seen as solvers for a relaxation of our QAP.
Appendix B Mutual k-NN and CKA as distortion metrics
In Sec. 4, we introduce the notation of distortion metrics in terms of a unimodal kernel function and , and a distance function , which is decomposable as
In this section, we show how Mutual k-NN [17], the centered kernel alignment (CKA) [22], and the Gromov-Wasserstein () distance [33] fit into this formulation.
[…]
which trivially satisfies Eq. 15 with and , being identity functions. This choice of kernels and the distance metric leads to the Mutual k-NN metric in our framework:
[…]
where . Similar to previous work [30], we use the linear kernel in this work. Our kernel matrices are
Models | CIFAR-10 (%) | CINIC-10 (%) |
Mutual k-NN | ||
CKA | ||
distance |
[…]
distance [33]. For the distance, we can choose the -norm as the kernel and the squared distance as the distortion metric:
[…]
The objective function after finding the optimal permutation matrix similar to Eq. 6 is then equivalent to the original distance comparing two metric-measure spaces [38].
Tab. 3 empirically compares these formulations in a small-scale data regime, using CIFAR-10 [23] and CINIC-10 [9] datasets introduced in the main text. We find that despite the wider adoption of Mutual k-NN and CKA metrics in previous work, the Gromov-Wasserstein distance leads to a higher matching accuracy on both datasets.
Appendix C Shuffled vision-language alignment
In this section, we elaborate on the setup for the shuffling experiment from Sec. 3. As we claimed in the main text (cf. Sec. 3), the observations are consistent accross all tested datasets and models, which we also report here.
Setup. Given aligned image and language representations, , and a shuffling level , we randomly choose elements that are randomly permuted. Every other element is kept in place. Afterwards, the distortion of this permutation is computed with Eq. 3 or Eq. 4 after the permutation. Note that we use the image embeddings here instead of the averaged object embeddings. For classification datasets, we take the same language embedding for all elements of that class. However, we have seen in preliminary experiments that the curve looks similar when considering the averaged object representations instead of the image representations. We plot equidistant shuffling levels at . Each level is based on random seeds to sample the subset and permutation.
Shuffling with other kernels, datasets, and models. In addition to Fig. 2, which uses the CocoCaptions dataset [6], and Mutual k-NN as the distortion metric, we show more combinations in Fig. 7. In this setting, Mutual k-NN is only meaningful for paired datasets because the k-nearest neighbors are ambiguous when language features are replicated. For the distance, the pairwise distances are also dependent on the dimensionality of the embedding spaces. Therefore, we standardize the distance to be in the range between zero and one for this shuffling experiment. We observe that the similarity (/ distortion) decreases (/ increases) strongly monotonically with more shuffling. This behavior is consistent for all considered datasets and metrics. This observation suggests that the pairwise relations are more similar between the semantically corresponding vision and language representations than between the non-semantic ones.
A note on randomly initialized networks. In Fig. 8, we present a plot of the experiment in Fig. 2 with the addition of randomly initialized ViT-H/14 [11] models, which were initiated with distinct random seeds. Additionally, we demonstrate the distortion of a representation based on the stacked pixel values, i.e., without any neural network. Furthermore, we show a zoomed-in version on the left-hand side to illustrate the behavior on a finer scale.
We observe that the alignment of randomly initialized networks exhibits the same monotonically decreasing behavior with increased amount of shuffling, as for pre-trained networks (despite the absolute alignment value being smaller). For images, this can be explained by the similarity of the pixel distribution within each semantic category (e.g., a green landscape for animal stock). These similarities appear more frequently for semantically affiliated classes and can dominate the pairwise distance encoding. To understand the behavior for randomly initialized networks, we plot the distribution of the Empirical Lipschitz constant in Fig. 9, defined by
for samples , and function . Intuitively, it measures the degree of distance distortion in the output space w.r.t. the input space for each sample pair in the dataset. Here, we use CocoCaptions and the same language model as in Fig. 8. We observe that most values are close to one. This implies that the distance after encoding remains approximately preserved. Nevertheless, the distances are slightly distorted, which explains why the absolute similarity in Fig. 8 is lower for random ViTs than for the pixel values. In summary, shuffling reduces alignment even for the distance in terms of pixel values. This then transfers to random ViTs because these distances are approximately preserved.
Appendix D On the factorized Hahn-Grant solver
In Sec. 4, we recap the Hahn-Grant solver [13] and introduce our factorized Hahn-Grant solver. Here, we provide more details on the implementation of the algorithms and show that both algorithms result in the same lower bounds. We also include an empirical study supporting the design choices behind our factorized Hahn-Grant solver: the factorization, faster LAP solvers, and finding primal solutions.
D.1 Proof of equivalence to Hahn-Grant solver
The main idea of the proof is that the lower bound, the leader, and all LAPs are the same for both algorithms. The main difference is that the dual vectors are stored in the tensors and instead of updating in-place.
[…]
Thus, both algorithms have an equivalent starting condition.
Next, given that Eq. 27 holds, we show that both values are changed in an equivalent way. Given the same leader matrix, Line 6-8 from Algorithm 1 change and leader in the same way as Lines 6-9 from Algorithm 2. Moreover, after Algorithm 1, Line 9 and Algorithm 2, Line 10, and are changed in the same way. Therefore, the sum is also preserved:
where the superscipt denotes the value before executing the line.
In Line 11 and Line 12 of Algorithm 1, the values from the complementary positions are redistributed to the current submatrix. Line 12 from Algorithm 2 also aggregates the sum of the complementary elements, albeit not by changing the cost matrices. Therefore, the sum of both elements remains the same. By definition of , it follows, that and the Hungarian matching produces the same values for both algorithms. As the next step, the objective value is added to the leader in Algorithm 1 by first adding it to in Line 14 and then to in Line 5 in the next iteration. In Algorithm 2, this value is directly added to . Finally, the dual variables are subtracted from the cost in Line 15 of Algorithm 1 and Line 16 and Line 17 of Algorithm 2. This preserves the sum:
Therefore, each iteration in both algorithms changes the costs in an equivalent way. As a result, the final bound and each solution to the LAPs are the same.
[…]
Therefore, we can subtract the minimal element from both matrices, apply the algorithm to the resulting non-negative matrices, and subtract the constant from Eq. 32 to the final objective value to retrieve the optimal objective value of the original problem.
D.2 Implementation details
We implement both Algorithm 2 and Algorithm 1 in Python using PyTorch [2]. The main computational bottleneck is the LAP solver. Therefore, we use a custom C++ algorithm for the forward-reverse auction [3] algorithm and for the Jonker-Volgenant algorithm [19, 31]. We stop the algorithm if the relative or absolute improvement of the dual bound is smaller than or the primal objective is within of the dual bound. Finally, for the auction algorithm, a larger relaxation usually leads to a faster runtime with the drawback of worse objectives and dual vectors. The Hahn-Grant algorithm also works for suboptimal dual vectors, but we observe that towards the end of the algorithm, better solutions in the LAPs are required. Therefore, we initialize relatively high in the beginning and multiply it with a factor of in every iteration.
Solver | Cost | Bound | ||
Hahn-Grant | -1.979877 | – | ||
+ factorized | -1.979875 | -2.089994 | ||
+ auction | -1.979871 | -1.981942 | ||
+ Jonker-Volgenant | -1.979878 | -1.980387 | ||
+ LAP solutions | -1.979914 | -1.980387 | ||
+ primal heuristic | -1.980015 | -1.980387 |
D.3 Ablations
Setup. We compare the original Hahn-Grant solver with our adaptation from Sec. 4: the factorization into matrices and , using the auction algorithm or the Jonker-Volgenant as faster LAP solvers (Line 6 and Line 13), evaluating the individual LAP primal solutions (Line 7 and Line 14), and using primal heuristics (Line 4). For these experiments, we generate two sets with vectors of dimensionality each, drawn element-wise from a standard normal distribution. These vectors are normalized, and the pairwise inner product is computed to produce two similarity matrices. We apply all variations of the algorithms with a time limit of two hours to these cost matrices and evaluate the quality of the primal solutions and of the bound. We repeat the experiment with random seeds and average the results. We also repeat the experiment with a set of random vectors and a time limit of one hour.
Results. Tab. 4 shows the result for each adaptation for size and Tab. 5 for size . We observe that the original Hahn-Grant solver reaches a non-trivial primal-dual bound for size but does not finish the first iteration within the two hour time limit for size . Even though factorization was introduced for improved memory, it also slightly speeds up the computation, leading to a smaller bound for both sizes. The small increase in the objective for size can be explained by the fact that the returned primal estimate is only the primal solution of the LAP for the leader. Therefore, the quality of this estimate can vary with every iteration. The auction and Jonker-Volgenant algorithms were introduced to improve the speed of the algorithm and, therefore, lead to smaller bounds. However, the Jonker-Volgenant algorithm leads to the tightest bounds empirically for large problems while the auction algorithm is slightly better on the size problem. The LAP solutions do not change the dual estimate but improve the primal solution by a significant margin. Finally, the primal heuristics further improve the primal solution. Since we are mostly interested in good primal solutions and measuring the quality of the solution, we use the algorithm with all the adaptations.
Solver | Cost | Bound | ||
Hahn-Grant | -1.950160 | -2.245823 | ||
+ factorized | -1.950166 | -1.951418 | ||
+ auction | -1.950197 | -1.950586 | ||
+ Jonker-Volgenant | -1.950197 | -1.950623 | ||
+ LAP solutions | -1.950268 | -1.950623 | ||
+ primal heuristic | -1.950360 | -1.950623 |
Appendix E Experimental details
In this section, we provide more details supplementing the experiments in Sec. 5, which were excluded from the main text due to space constraints.
Our method only assumes pre-computed embeddings (or already pre-computed similarity matrices). Therefore, we are not limited to specific architectures or pre-training strategies. Our goal is to choose a variety of different architectures, pre-trainings, and model sizes for both modalities.
Vision models. We consider self-supervised, fully-supervised and vision-language supervised models with convolutional and vision transformer architectures of different capacities. For self-supervised methods, we consider different models from DINO [3] trained on ImageNet-1k [44] and models from DINOv2 [36] trained on LVD-142M. For supervised models, DeiT [48] and ConvNeXt [27] pre-trained on ImageNet-1k, ImageNet-22k [44] and a a combination of both are used. Furthermore, we choose CLIP [42] as our vision-language model with both ResNet and ViT backbones. We use the code and models from the official repository except for ConvNeXt, where we use pre-trained models from the timm library. In total, we use vision models.
Language models. Similar to the vision models, we consider different pre-trainings and network sizes. In particular, in addition the the CLIP [42] text models, contrastive learning, question-answer models, and average word embeddings are considered. We use the official CLIP repository for all CLIP text models and the SentenceTransformers [43] library for all other models. In total, we use language models.
Datasets. We evaluate our experiments on CIFAR-10 [23], CINIC-10 [9], CIFAR-100 [23], and ImageNet-100 [44]. For ImageNet-100, we use the validation split and for all other datasets, we use the test split. We choose the same class names and prompts as ASIF [35] whenever available, and follow the same preprocessing as CLIP otherwise. For ImageNet-100, we observed an improved performance by encoding each class as “name: definition“ using the WordNet definitions and more descriptive class names.
General setup. We implement all experiments in Python with PyTorch [2]. In general, we pre-compute the embeddings for each vision and language model and every dataset using PyTorch Lightning [4]. We do not use GPUs except for precomputing the embeddings. After computing the image-wise embeddings, we normalize them and average them for every class, see Eq. 1. The resulting object-wise embedding is again normalized. To evaluate the impact of small changes in the embeddings (and pairwise similarities), we only take a random subset of the image representations to compute the average. In particular, we use a half of the embeddings drawn uniformly for every random seed. For the language embeddings, we take the average over the embeddings for all different prompts.
Small-scale matching. We use random seeds and all models on CIFAR-10 and CINIC-10 for small-scale matching. In each experiment, we enumerate all permutations and compute all costs explicitly, returning the permutation with the minimal cost. The comparison of all combinations of vision and language models is given in Sec. F.1.
Larger-scale matching. For larger-scale matching, we use different subsets of classes using the optimization problem introduced in Sec. 4.2 solved with Gurobi [12]. For each problem size , we use the best subset of classes, each with random seeds for computing the object-level vision embeddings. We use our factorized Hahn-Grant solver and a time limit of one hour for each matching problem. The models are DINOv2 ViT-G/14, CLIP ViT-L/14@336px, and the distilled DeiT-B/16@384px with all-mpnet-base-v2.
Solver comparison. We evaluate the solvers both on small-scale and larger-scale matching. For small-scale matching, we use random seeds for DINOv2 ViT-G/14, CLIP ViT-L/14@336px, and the distilled DeiT-B/16@384px with all-mpnet-base-v2 and All-Roberta-large-v1 on CIFAR-10 and CINIC-10. We present the results for all combinations in F.3. The larger-scale benchmark follows the setting of larger-scale matching, but only considers one out of the ten subsets and one random seed for CIFAR-100 using CLIP CLIP ViT-L/14@336px and all-mpnet-base-v2 with a time limit of 1.5 hours.
Unsupervised classifier. For unsupervised classification, we use K-Means [28] to cluster image embeddings into prototype (object) embeddings. We use K-Means++ [5] from Scikit-learn [1] with initializations. The cluster centers are then matched to the language embeddings using our factorized Hahn-Grant solver. Similar to the previous experiments, we only use a random 50% subset of the vision embeddings and evaluate the method for random seeds. We report the results for DINOv2 ViT-G/14, CLIP ViT-L/14@336px, and the distilled DeiT-B/16@384px with all-mpnet-base-v2 and All-Roberta-large-v1 on CIFAR-10.
Solver | Accuracy (%) | Matching Accuracy (%) |
Random | ±10.3 | ±11.6 |
LocalCKA | ±7.0 | ±7.7 |
OT | ±9.5 | ±9.4 |
FAQ | ±8.7 | ±9.9 |
MPOpt | ±21.4 | ±23.9 |
Ours | ±3.4 | ±3.7 |
GT | ±0.6 | ±0.0 |
Appendix F Evaluation results
Following up on Sec. 5, we report the results for all vision and language models in the small-scale matching setting (cf. Sec. F.1) and benchmark of the different solvers using multiple datasets and models (cf. Sec. F.2 and Sec. F.3).
F.1 Comparison of vision and language models
In this section, we report the results on the small-scale matching, spanning all vision and language models in our study. The results show, in particular, that DINOv2 outperforms the other models on both datasets and that the model size is less important than the pre-training method.
In Fig. 11, we show the matching accuracy (top) and Mutual k-NN with (bottom) of each vision model in combination with All-Roberta-large-v1 (left) and all-mpnet-base-v2 (right) for CIFAR-10 (left) and CINIC-10 (right). The lines show the trend for increasing model sizes that are fitted to the models of varying capacity for each model class (different colors). The model size corresponds to the number of parameters. First, we observe that for both datasets and language models, DINOv2 yields the highest matching accuracy. Furthermore, for every model class, there is no a clear propensity of larger models to perform better. As this seems to contradict the platonic representation hypothesis Huh et al. [17], we report the Mutual k-NN in the bottom plots. We observe that in line with our conclusions, there is still no clear improvement of the Mutual k-NN metric w.r.t. an increasing model capacity. This implies that larger models do not necessarily lead to stronger alignment with language – at least on CIFAR-10 and CINIC-10. Considering the observations by Huh et al. [17], this suggests that scaling the models could lead to a better alignment on the Wit dataset [6], even though it may not be sufficient to improve the alignment on every other dataset. We report the individual matching accuracy of each combination of vision and language model for CIFAR-10 in Fig. 12 and for CINIC-10 in Fig. 13.
F.2 Unsupervised classification: solver comparison
We also compare the solvers on the unsupervised classification setting in Tab. 6. In addition to the classification accuracy, we also report the matching accuracy. Given a clustering and the ground truth labels, we compute the optimal ground truth matching. The matching accuracy evaluates how well our unsupervised matching coincides with the ground truth matching. Finally, we report the performance when the ground truth matching is used instead of our unsupervised matching to evaluate the quality of the clustering.
Similar to the unsupervised matching experiments, we observe that our solver outperforms the other solvers in terms of both classification accuracy and matching accuracy. This means that our solver finds matches that agree well with the ground truth matches. We also observe that the k-means clustering is not perfect, and approximately of the images are clustered in the wrong clusters. This hurts our clustering in two ways. First, since only whole clusters are assigned to classes, misclassified images will remain in the respective (wrong) cluster. Second, incorrect samples distort the centroids, which in turn can further affect the pairwise distances, leading to a worse matching. However, even with imperfect centroids, our matching finds non-trivial matches and leads to the first non-trivial unsupervised classification, which was not possible using previous local solvers.
F.3 Comparison of solvers on small-scale problems
Similar to Tab. 1, we benchmark different solvers for all combinations in Tab. 7. Our factorized Hahn-Grant solver and Gurobi [12] find the global optimum for all problems. For most combinations, these solvers are also the only solvers finding the global optimum. MPOpt [18] also finds the global optimum for some problems (e.g., CIFAR-10 with DINOv2) but fails for other problems (e.g., CIFAR-10 with CLIP). For all but two experiments (CIFAR-10 with DeiT and all-mpnet-base-v2 and CINIC-10 with DeiT and All-Roberta-large-v1), the global optimum also leads to the best matching accuracy, which underlines the importance of the Gromov-Wasserstein distance as a proxy measure for blind matching. Finally, the matching accuracy is close to random for most solvers (LocalCKA [30], OT [38], and FAQ [49]). This shows that a global solver is crucial to obtain non-trivial results.
All-Roberta-large-v1 | all-mpnet-base-v2 | |||||||
Accuracy (%) | Cost | Global? (%) | Accuracy (%) | Cost | Global? (%) | |||
CIFAR-10 | CLIP | Random | ±10.4 | ±0.32 | ±0.0 | ±10.4 | ±0.29 | ±0.0 |
LocalCKA [30] | ±8.87 | ±0.33 | ±0.0 | ±11.64 | ±0.31 | ±0.0 | ||
OT [38] | ±3.08 | ±0.03 | ±0.0 | ±0.0 | ±0.01 | ±0.0 | ||
FAQ [49] | ±3.66 | ±0.06 | ±0.0 | ±8.89 | ±0.02 | ±41.04 | ||
MPOpt [18] | ±0.0 | ±0.01 | ±0.0 | ±0.0 | ±0.01 | ±0.0 | ||
Gurobi [12] | ±10.05 | ±0.01 | ±0.0 | ±4.44 | ±0.01 | ±0.0 | ||
Ours | ±10.05 | ±0.01 | ±0.0 | ±4.44 | ±0.01 | ±0.0 | ||
DeiT | Random | ±10.4 | ±0.29 | ±0.0 | ±10.4 | ±0.27 | ±0.0 | |
LocalCKA [30] | ±9.18 | ±0.2 | ±0.0 | ±6.81 | ±0.22 | ±0.0 | ||
OT [38] | ±0.0 | ±0.01 | ±0.0 | ±3.08 | ±0.01 | ±47.02 | ||
FAQ [49] | ±10.7 | ±0.13 | ±0.0 | ±14.9 | ±0.01 | ±0.0 | ||
MPOpt [18] | ±0.0 | ±0.02 | ±0.0 | ±0.0 | ±0.01 | ±0.0 | ||
Gurobi [12] | ±4.7 | ±0.01 | ±0.0 | ±6.81 | ±0.0 | ±0.0 | ||
Ours | ±4.7 | ±0.01 | ±0.0 | ±6.81 | ±0.0 | ±0.0 | ||
DINOv2 | Random | ±10.4 | ±0.27 | ±0.0 | ±10.4 | ±0.28 | ±0.0 | |
LocalCKA [30] | ±11.42 | ±0.31 | ±0.0 | ±9.1 | ±0.24 | ±0.0 | ||
OT [38] | ±26.25 | ±0.44 | ±0.0 | ±24.15 | ±0.54 | ±22.36 | ||
FAQ [49] | ±35.3 | ±0.18 | ±22.36 | ±23.82 | ±0.22 | ±0.0 | ||
MPOpt [18] | ±0.0 | ±0.01 | ±0.0 | ±17.65 | ±0.01 | ±0.0 | ||
Gurobi [12] | ±0.0 | ±0.01 | ±0.0 | ±17.65 | ±0.01 | ±0.0 | ||
Ours | ±0.0 | ±0.01 | ±0.0 | ±17.65 | ±0.01 | ±0.0 | ||
CINIC-10 | CLIP | Random | ±10.4 | ±0.31 | ±0.0 | ±10.4 | ±0.3 | ±0.0 |
LocalCKA [30] | ±8.87 | ±0.25 | ±0.0 | ±8.58 | ±0.27 | ±0.0 | ||
OT [38] | ±3.66 | ±0.13 | ±0.0 | ±3.66 | ±0.11 | ±0.0 | ||
FAQ [49] | ±5.5 | ±0.02 | ±0.0 | ±3.08 | ±0.08 | ±0.0 | ||
MPOpt [18] | ±2.24 | ±0.04 | ±22.36 | ±12.1 | ±0.02 | ±51.3 | ||
Gurobi [12] | ±0.0 | ±0.01 | ±0.0 | ±0.0 | ±0.02 | ±0.0 | ||
Ours | ±0.0 | ±0.01 | ±0.0 | ±0.0 | ±0.02 | ±0.0 | ||
DeiT | Random | ±10.4 | ±0.25 | ±0.0 | ±10.4 | ±0.25 | ±0.0 | |
LocalCKA [30] | ±7.88 | ±0.17 | ±0.0 | ±8.75 | ±0.22 | ±0.0 | ||
OT [38] | ±6.57 | ±0.02 | ±0.0 | ±4.7 | ±0.05 | ±0.0 | ||
FAQ [49] | ±17.01 | ±0.07 | ±0.0 | ±3.08 | ±0.04 | ±0.0 | ||
MPOpt [18] | ±4.89 | ±0.01 | ±51.04 | ±6.16 | ±0.01 | ±0.0 | ||
Gurobi [12] | ±3.08 | ±0.01 | ±0.0 | ±6.16 | ±0.01 | ±0.0 | ||
Ours | ±3.08 | ±0.01 | ±0.0 | ±6.16 | ±0.01 | ±0.0 | ||
DINOv2 | Random | ±10.4 | ±0.23 | ±0.0 | ±10.4 | ±0.23 | ±0.0 | |
LocalCKA [30] | ±9.45 | ±0.14 | ±0.0 | ±7.33 | ±0.21 | ±0.0 | ||
OT [38] | ±15.72 | ±0.06 | ±48.94 | ±5.03 | ±0.22 | ±0.0 | ||
FAQ [49] | ±13.33 | ±0.06 | ±0.0 | ±10.21 | ±0.12 | ±0.0 | ||
MPOpt [18] | ±0.0 | ±0.02 | ±22.36 | ±13.8 | ±0.03 | ±30.78 | ||
Gurobi [12] | ±0.0 | ±0.02 | ±0.0 | ±7.33 | ±0.02 | ±0.0 | ||
Ours | ±0.0 | ±0.02 | ±0.0 | ±7.33 | ±0.02 | ±0.0 |
Appendix G Optimal transport as QAP solver
We show that optimal transport with uniform source and target probability distributions is equivalent to relaxing the QAP with doubly stochastic matrices. In our experiments, we use this equivalence to compare to the solutions from optimal transport solvers.
[…]
Here, we use the fact that is a valid transport matrix as and . The other direction can be derived in a similar way by dividing the doubly stochastic matrix by .